OpenStudy (anonymous):

write the slope intercept form of the equation of the line that passes through (4,2) with a slope of 3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(4\quad ,&2) \end{array} \\\quad \\ slope = m= 3 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)